When 34.2 grams of strontium reacts with excess molecular oxygen, how many grams of the product will form?
Answer #1

Molar mass of Sr = 87.62 g/mol

mass of Sr = 34.2 g

mol of Sr = (mass)/(molar mass)

= 34.2/87.62

= 0.3903 mol

Balanced chemical equation is:

2 Sr + O2 —> 2 SrO

According to balanced equation

mol of SrO formed = moles of Sr

= 0.3903 mol

Molar mass of SrO,

MM = 1*MM(Sr) + 1*MM(O)

= 1*87.62 + 1*16.0

= 103.62 g/mol

mass of SrO = number of mol * molar mass

= 0.3903*1.036*10^2

= 40.45 g

Answer: 40.5 g